OpenStudy (anonymous):

How many unique triangles can be made where one angle measures 98° and another angle is an obtuse angle? A) 1 B) 2 C) More then 2 D) None

OpenStudy (gorv):

u know what a obtuse angle means ??

OpenStudy (gorv):

obtuse angle are the angle > 90 degree

OpenStudy (gorv):

in one triangle how many angle greater than 90 degree are possible

OpenStudy (gorv):

??

OpenStudy (gorv):

@Kittymeowb

OpenStudy (anonymous):

1?

OpenStudy (gorv):

yeah ..and what your Q say??

OpenStudy (anonymous):

one angle is 98 and the other obtuse?

OpenStudy (gorv):

so is it possible to have 2 angle greater than 90???

OpenStudy (anonymous):

yes?

OpenStudy (gorv):

how ??

OpenStudy (gorv):

sum of all the angle of triangle = 180

OpenStudy (gorv):

98+91+x=180 189+x=180 x=180-189=-9 x=-9degree

OpenStudy (gorv):

it is not possible

OpenStudy (anonymous):

o-oh, but how does the angles matter on how many unique triangles i can make?

OpenStudy (anonymous):

Obtuse angle means angle which is greater than \(90\).. In a triangle only one angle can be obtuse, so one angle can be greater than 90..

OpenStudy (anonymous):

One angle is given to you, that is greater than 90 too, one angle that is given to you is already obtuse..

OpenStudy (anonymous):

So, can you make one more obtuse angle in your triangle when you are given one obtuse angle already?

OpenStudy (anonymous):

so does that mean the answer would be none? since you said only one can be obtuse?

OpenStudy (anonymous):

Exactly...

OpenStudy (anonymous):

you can make only one obtuse angle in any triangle and it is given to you already, so further you cannot make any other obtuse angle in the triangle.. So, \(\text{None}\) is the best answer to your question.. :)

OpenStudy (anonymous):

\(\text{This is it...}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you are welcome dear.. :)